What charge will be required to liberate t moles of hydrogen gas if dilute HCl is electrolysed using inert electrodes for 5 minutes ? Calculate the current required ?


1
Expert's answer
2021-05-03T07:35:30-0400

1 mol of H2 will require 2×1=2 moles of electrons or 2 faradays.

No of Faraday=Charge/96500

2=(I×5×60)/96500

I=(96500×2)/(5×60)

I=64.3mA

Hence the charge is 64.3mA